X Band Gregorian Dome Receiver Log

as of Saturday, 16-Dec-2017 22:07:59 AST

Go to Most Recent Entry

NameEmailDateComment
ANS anolla@naic.edu 12/21/2001 The X band recvr was installed on 12/19/2001. Found that the bias box was not wired to monitor the cryo temp downstairs ( no microtech connector on it).
ANS anolla@naic.edu 12/21/2001 Try the bias box of the 3 ghz recvr , but didn't work ,something wrong with it. Also the biases for the rf amps are the following: Channel A stage 1 2 3 Vd (V) 1.48 2.0 2.5 Id(ma) 6.1 5.2 5.1 Channel B Vd (V) 2.0 2.51 2.53 Id(ma) 6.2 7.5 10.0
ANS anolla@naic.edu 1/21/2002 Channel A post amp bad , post chassis taken down on 1/18/2002.There are not spares for it.
ANS anolla@naic.edu 1/23/2002 Ganesh found a spare Milliwave amp at Ithaca lab on one of his trips. That's the one to be used to replace the burned one.
ANS anolla@naic.edu 1/24/2002 System installed back, Tsys on chann A is 39 ēk, and Tsys on chann B is 48 ēk.
ANS anolla@naic.edu 2/1/2002 Post amps chassis taken down in order to install extra post amps on it.
ANS anolla@naic.edu 2/6/2002 System focal point was adjusted by .44". Shims are now 2.714".
ANS anolla@naic.edu 2/11/2002 New horn heater power supply installed.
ANS anolla@naic.edu 2/18/2002 System cryocomp running a third cryopump , the NLB one. It's cryocomp vent fan failed.
ANS anolla@naic.edu 3/4/2002 Channel A was with a Tsys of 100 ēk instead the normal 43 ēk , chann B was normal. Ganesh found that the exiter LED's were off. For some reason chann B didn't care if this LED is on or off, but not chann A.Also JR took the bias readings and are : Chann Vd1 Id1 Vd2 Id2 Vd3 Id3 A 1.5 vd 6.2 ma 2.0 vd 5.2ma 2.5 vd 5.1ma B 2.02 vd 6.2ma 2.52 vd 7.5ma 2.53 vd 10ma
ANS anolla@naic.edu 5/13/2002 System hot due to cryocomp failure ,power dip ,it takes to long to cool down because there are three receivers in same compressor.Dewar valve left open.
ANS anolla@naic.edu 5/14/2002 Receiver Dewar valve closed. 20.1 / 95.7 /VP 6 .
ANS anolla@naic.edu 5/17/2002 System was hot due to cryocomp failure because a power dip.Was cooled.
ANS anolla@naic.edu 5/28/2002 The protection shutter at the horn was left on it by accident since the last SB xmitter run on 5/24/2002. This morning during a T sys Norberto Despiau noticed that the T sys for these system was above 100ēk .JR & IC double check it and found that system was still protected.
ANS anolla@naic.edu 7/31/2002 On 6/21/2002 cryocomp 009 failed (cooling fan stop working) , it was replaced with 011.
LSWlwray@naic.eduFriday, August 16, 2002 16:08:25Power glitch caused 3-receiver compressor (LBN, SBN, X) to turn off, 1.5 hours later it was noticed. The vacuum pump is running on this receiver, it will not be cooled until Monday.
ansanolla@naic.eduMonday, September 9, 2002 15:52:09The Dewar warmed up on 9/8/2002 due to a power dip that shut off the compressor.Nobody noticed it until several hours later. So we started its cool down this morning,
ansanolla@naic.eduTuesday, October 8, 2002 09:36:15System down , bias at lab to check it. It has something bad when we try it with the new cryo temp reader.
ansanolla@naic.eduWednesday, November 13, 2002 08:59:47The vacuum was lost at the Dewar. Was pumped and return to 22 & 110ēk
LSWlwray@naic.eduWednesday, November 27, 2002 20:49:21Monday Nov 25 - phil and slava's experiment showed a 833 ms square wave in their ch B data. this is the period of the crosshead (1.2 Hz).
ansanolla@naic.eduMonday, December 9, 2002 12:47:40System has cryogenics problem.It was cooled down , but its cryo temps were to high . 31 & 67 ēk . In order to check if it was due running 3 recvs with one comp making the XB marginal we switched to the SBW comp that was free ,but cryo tems still to high.So is something else.
ansanolla@naic edu.Thursday, December 12, 2002 13:37:33Cryopump replaced again , the ED 013 that was installed as replacement of ED 007 didn't cooled the Dewar as suppose to do.It was 30 & 66 deg k instead 15 & 50 deg K. Now cryopump ED 014 is at it.
ansanolla@naic edu.Friday, December 13, 2002 09:18:03Cryotemps are 23.2ēk & 57.9ēk after replacing cryopump for second time in a rool.
ansanolla@naic edu.Monday, December 16, 2002 12:39:49The VP is reading 29 , 16ē stage 24 and 70ēstage 55.7.Dewar not sweating.
ansanolla@naic edu.Friday, January 17, 2003 13:51:09Dewar cryogenics transfered from cryocomp 011 to cryocomp 009 (from trio recvrs Dewar to SBH Dewar). Due that the load ( 3 ea 350 cryopumps) of the trio receivers was not letting the NLB get cool properly.Now in cryocomp 009 there are a 350 & 1020 cryopumps.
ansanolla@naic.eduTuesday, January 21, 2003 14:43:24Cryo lines 19/20 were replaced with cryo lines 17/18. Lines 19/20 been finished by Jose Vives.
ansanolla@naic edu.Thursday, January 23, 2003 16:00:22System was shut off due to contamination of cryo system due to new cryo lines.
ansanolla@naic edu.Thursday, February 20, 2003 14:56:39Cryopump replaced by Coco. Also system was cooled down ,cryo temps are 18.9 ēk & 45.7 ēk.
ansanolla@naic edu.Monday, March 3, 2003 16:26:42Changed attenuators after post amp from: 6/6 dB (A/B) to 14/14 dB, rebalancing levels after upstairs RF switch change.
lswlwray@naic.eduThursday, March 13, 2003 17:07:46Put protective hat on horn for the SB radar run.
lswlwray@naic.eduWednesday, March 26, 2003 16:25:34Moved XB from XB/SBH compressor (008) to SBN/LBN/XB (011).
ansanolla@naic.eduThursday, June 26, 2003 15:18:06The bias chassis was removed from receiver in order to add an splitter , it was not posible , so it was installed back to do it in other time.
ansanolla@naic.eduTuesday, July 1, 2003 15:54:22The XB chassis down in order to add the splitter for Tim Hankings.
ansanolla@naic.eduWednesday, July 23, 2003 09:08:20System warmed up due to power dip. Operator tried to restart cryo comp but was not able. Later we found that a relay in the restart ckt was not working.
lswlwray@naic.eduWednesday, September 3, 2003 15:00Changed horn from cylindrical one to wide-flare. Also cleaned cal cables (some dirt was in there). Noticed reduced tsys afterwards.
lswlwray@naic.eduFriday, September 5, 2003 15:00Added shutter to new horn.
ansanolla@naic.eduMonday, September 8, 2003 15:21:10Channel A had an unstability problem . It was due to a loose connector inside the RF cabinet.The connector at the feed thru was not tighten properly.
ansanolla@naic.eduFriday, September 19, 2003 15:07:42System down for maint.Until further notice.
ansanolla@naic.eduThursday, October 9, 2003 12:48:31Receiver was installed back after a series of test.It's been cooled.
ansanolla@naic.eduMonday, November 10, 2003 13:20:29System warmed up on Saturday due to cryocomp (ED 011) failure.Also JRV replacesd the cryopump ( It was ED ,now is ED 007.) System still warmed due to cryopump replacement)
ansanolla@naic.eduTuesday, November 18, 2003 11:53:49System was cooled down.It was done alone , so system cryo temps were 19.5 & 59.7 and vpg was 6. But after adding the NL & NS the Dewar started misbehaving. It cryo temps increased to 27.8 & 65.7 and it's vpg reached the 60's.
ansanolla@naic.eduTuesday, December 16, 2003 14:21:27System down (chassis down for modifications,rf & bias).
ansanolla@naic.eduWednesday, December 17, 2003 13:35:57System back to normal.Due to 430 mhz xmitter failure all Dome receivers back to normal.
ansanolla@naic.eduWednesday, January 7, 2004 15:03:53The XB receiver will be rellocated at the former NLB location at the rotary floor.
lswlwray@naic.eduFriday, January 30, 2004 12:48:34Problem with cryogenic lines serving the SBN and XB receivers. Leak somewhere. Receivers are warm until further notice.
lswlwray@naic.eduMonday, February 2, 2004 15:56:36Problems with cryogenic lines avoided by moving the lines to old#13 from receiver room to transmitter room and new#17 from transmitter room to compressor area. Receiver is cooling.
ansanolla@naic.eduFriday, February 13, 2004 16:13:52System down for modification.Going to replace rf amps at the Dewar.
ansanolla@naic.eduWednesday, March 3, 2004 10:13:10The amps been replaced , post amps by passed due to gain on new rf ones. Also the Micronetics noise cal diodes were replaced by Noise Com ones.
ansanolla@naic.eduMonday, March 15, 2004 16:11:40System installed back after replacing rf amps. Also cal diodes were replaced .The cryo temp reading ckt is not working , have to be checked.
ans anolla@naic.eduMonday, April 5, 2004 11:24:00XB was taken down since 3/29/2004 in order to fix the remote cryo temp reading. Today was hooked up back.
ansanolla@naic.eduTuesday, May 25, 2004 15:45:21The horn seal cover damage . Also the waveguide on horn is corroded , need's service.
lswlwray@naic.eduWednesday, June 2, 2004 16:12:57This morning, the compressor that runs the SBN and XB failed (flappers). The XB and SBN receivers warmed up. They were put onto the SBW's compressor (the receiver is down due to cryo problems) and cooled. Both receivers are ready as of 4pm.
lswlwray@naic.eduMonday, June 7, 2004 13:20:21Bandpass filter (7.8 - 10.2 GHz) removed from postamp chassis today, in order to observe at 10.45 GHz tonight.
lswlwray@naic.eduTuesday, June 22, 2004 14:01:04Replaced bandpass filters today. Tsys dropped down to normal at ~30K each channel (excess noise allowed by filter's absence was mixing in.
ansanolla@naic.eduFriday, July 2, 2004 11:36:12System down due to power shut down (main power line shorted at the rotary joint).
ansanolla@naic.eduTuesday, July 6, 2004 10:48:43System been cooled.Started this morning.
ansanolla@naic.eduTuesday, July 6, 2004 15:37:14System back to normal , it was down since 7/2/2004 due to a main power line shorted out at the rotary joint.
ansanolla@naic.eduThursday, July 15, 2004 16:22:08System cryocompressor swap from ED 012 to ED 008 due to vibration on it.
ansanolla@naic.eduTuesday, July 27, 2004 13:38:41Due to power dip cryocomp ED 008 on 7/25/2004 at 13:35 stop working .It's cooling fan quit working overheating the unit.Spare cryocomp ED 12 attached to system , but there was a leaky hose ( return one) in these cryocomp making the system to warm up again.So for these reason SBN & XB were out of service.
ansanolla@naic.eduTuesday, July 27, 2004 13:47:00System been cooled.Cryocomp ED 009 on it
ansanolla@naic.eduFriday, July 30, 2004 11:42:33Cryopump got stuck due to contamination on 7/29/2004. It was purged & contamination ran away.Thanks Lord.
ansanoll@naic.eduTuesday, August 10, 2004 14:29:42XB cryo will be shut off in order to replace it's cryopump , is highly contaminated.
ansanolla@naic.eduWednesday, September 29, 2004 14:37:39Cryocomp replaced , it was ED 009 and was taken down for modification. Instead cryocomp ED 008 was installed.
ansanolla@naic.eduMonday, October 11, 2004 15:57:55System was down due to power failure on Saturday , was cooled down by BG and JRV.
ansanolla@naic.eduTuesday, November 23, 2004 16:18:36There it was a signal at 9.18 Ghz that was produced by an oscillation in the SBL chann A receiver. Bias was adjusted in the SBL and the oscillation was killed.
ansanolla@naic.eduWednesday, November 24, 2004 09:43:35System warmed up in order to replace the crosshead.
ansanolla@naic.eduMonday, November 29, 2004 13:32:27Cryohead replaced. It was ED 007 , is ED 015. System been cooled.
ansanolla@naic.eduTuesday, November 30, 2004 14:32:33The calibration system was not working properly in all the receivers at the Dome . It was due that the post amp chassis for the CBH and XB were off. These was loading the other recvrs calibrators.
ansanolla@naic.eduTuesday, January 4, 2005 15:57:20System warmed up due to failure of the SBN partner cryocomp.It was cooled again after been purged.
ansanolla@naic.eduWednesday, January 5, 2005 10:57:28XB cryohead lock up after been purged 2 times. So cryo system tured off until further notice.
ansanolla@naic.eduTuesday, January 11, 2005 13:42:15Cryocomp & lines were purged , contamination on them.Crosshead needs replacement.To be done tomorrow.
ansanolla@naic.eduWednesday, January 19, 2005 16:09:42Cryohead was replaced due to contamination.Crosshead motor shaft broken in two.Cooled down and it is ok now.
ansanolla@naic.eduMonday, January 31, 2005 11:30:20Noise injection board replaced.
ansanolla@naic.eduMonday, March 14, 2005 09:40:18Bias chassis at lab to replace the Id meter , was getting stock.Also power supplies will be swap (+/- 15vdc , add +5 vdc).
crioscrios@naic.eduFriday, June 24, 2005 13:25:14Tuesday June 21 the Shutter was fixed. The motor was replaced with a new one. System checked and working well.
ansanolla@naic.eduFriday, August 26, 2005 13:08:30Due to power failure on 8/25/2005 system warmed up.
ansanolla@naic.eduTuesday, August 30, 2005 15:46:02System cooled. Back to normal operation.
ansanolla@naic.eduFriday, September 9, 2005 13:55:32Shutter not working properly , its set screw was loose.Due to dual beam run shutter put on place & motor driver removed.
ansanolla@naic.eduTuesday, September 13, 2005 10:15:25Carlitos Rios fixed & check the shutter, working normaly.At least shutter mechanical operation should be check once in a week in all of them.
ansanolla@naic.eduTuesday, September 13, 2005 15:47:29There was not signal coming from upstairs due to the 2 ghz converter failure.Symptom is that the L.O. amplifier (SMC-10)is dead.Working on it.Signal diverted thru 10 Ghz converter.
ansanolla@naic.eduWednesday, September 14, 2005 13:46:37L.O. amp SMC-10 was dead ,replaced with the spare one & is ok now. Also the coax cable at the IF out inside the converter was bad (broken).
ansanolla@naic.eduWednesday, September 14, 2005 14:46:46Forgot to mention that the problem with the receiver was that the Low frequency converter (UG801D09) chassis had the LO amp (Miteq SMC-10) dead at the IF of 1.5 Ghz. Also a 3 db pad removed from the input of these LO amp , it had a 13 db att at its input now is a 10 db attenuation.
ansanolla@naic.eduThursday, September 15, 2005 13:23:34The shutter cover get loose again , set screw was set this time with Locktite .
ansanolla@naic.eduMonday, October 17, 2005 11:43:21System cryo head turned off accidentally by JG kids. been pumped.
ansanolla@naic.eduTuesday, October 18, 2005 12:45:00Dewar valve closed.System ok.
ansanolla@naic.eduTuesday, October 18, 2005 12:52:49New protective cap added to the on/off sw of the cryohead driver chassis. It was turned off accidentally yesterday by the Joe Greene kids.
ansanolla@naic.eduThursday, October 20, 2005 13:13:57A new safety divice added to the crosshead driver power sw in order to avoid to turn it off accidentally.
ansanollaTuesday, November 29, 2005 10:49:20Receiver cold.Take a look unto Dewar vac next maint.
ansanolla@naic.eduWednesday, November 30, 2005 13:40:07Receiver back to normal.
ansanolla@naic.eduWednesday, December 7, 2005 13:53:43The receiver Dewar is working ok since we vacuum it & cooled it.Why does it lost its vacuum? Tech answer ,who knows!!!!!!!
ansanolla@naic.eduThursday, December 15, 2005 13:33:29The remote cryo reading was not working properly.Found broken cable at the output of the remote reading connector on the bias chassis.We try two spare mux boards that were upstairs and didn't work , the third one did.
ansanolla@naic.eduMonday, December 19, 2005 10:59:14The vpg was reading 15 un's. So we started vacuum the Dewar.
ansanolla@naic.eduWednesday, December 21, 2005 11:46:05The Dewar valve was closed.Try to pinpoint Dewar vacuum leak but found that the vacuum gauge was not going lower than 50 un's.Receiver should be taken down to check the vac leak and to replaced the vacuum gauge.
ansanolla@naic.eduThursday, December 22, 2005 10:31:20System down for repairs.Bad vacuum gauge plus Dewar leak.
ansanolla@naic.eduWednesday, January 11, 2006 11:49:39The system installed back after vacuum problems with Dewar.Been cooled.
ansanolla@naic.eduThursday, January 12, 2006 09:00:48System cooled.
ansanolla@naic.eduMonday, February 27, 2006 11:08:56The 16ē stage reading bad , reading 1.3ēk localy. Remote reading with a wrong reading.
ansanolla@naic.eduMonday, July 31, 2006 13:10:43Check the problem with the 16 deg stage reading.Using the JRV gadget system was ok , have to check the bias cable for an intermittance or is coming from the Dewar.
ansanolla@naic.eduMonday, November 13, 2006 13:57:50The system was warming up.CR & JRV were upstairs checking what was going on.They started vacuum on Dewar . CR suspect tha cryohead is giving problems , so an eye on it.Receiver on schedule for 11/28/2006.
ansanolla@naic.eduTuesday, November 14, 2006 14:41:52System was not holding its 16 deg cryo temp.Suspect of the 16 deg seal.Being warmed up for a purge.Down until further notice.Operations notified.
ansanolla@naic.eduThursday, November 16, 2006 13:38:29System being cooled after purged , is suspected that the 16 deg seal is not holding.
ansanolla@naic.eduFriday, November 17, 2006 11:24:36System was cooled but the 16 deg stage still reading 36 K deg.So it will need to replace the cryo head , 16 deg seal bad.System being warmed.
ansanolla@naic.eduWednesday, November 22, 2006 16:13:50System being cooled.Vac pump on it.Cryo head replaced.
ansanolla@naic.eduMonday, November 27, 2006 13:03:38System didn't reach the 16 deg k at the 16 deg stage , bad seal on that stage? We will have to take a look on it.
ansanolla@naic.eduWednesday, November 29, 2006 14:18:47System was warmed up in order to check the problem with the 16 deg stage.It was found that the 16 deg seal was smash , was replaced and system being cooled.
ansanolla@naic.eduFriday, December 1, 2006 15:43:10Cryo head repleaced (is ED025 ) , system started a cool down at 14:30 AST today. New 16 deg stage seal on it.
ansanolla@naic.eduTuesday, December 5, 2006 15:35:22System has being suffering a lot of cryogenics problems.Last one was in its cool down on Friday 1,2006.The cryo head used to replace a bad one on it started giving leaking problems causing a major brakedown that add the SBN receiver cryo system onto the problem.Cryo comp and He lines are highly contaminated with oil & water.
ansanolla@naic.eduThursday, December 7, 2006 14:33:36System being cooled , cryo head replaced (is ED 020).Using the CBH cryo comp & lines ( cryo comp ED 010 & lines 21/22).
ansanolla@naic.eduFriday, December 8, 2006 11:01:24System back to normal. Cryo comp is ED 010 & cryo lines are 21 & 22.
crvcrios@naic.eduTuesday, December 12, 2006 13:23:52System now using SBH cryo Comp. and lines.(ed009, lines:17-18)
CRcrios@naic.eduMonday, April 16, 2007 10:03:50System taken down.
CRVThursday, January 3, 2008 09:31:11Vacuum pump connected to dewar ande pumping. Dic.13th,2007
CRVThursday, January 3, 2008 09:32:51Dewar cooldown Started on Dic.13th,2007 in the afternoon. Cold by DIc.13th during night time.
CRVThursday, January 3, 2008 09:34:14Dewar warmed out because the cross head of SBN broke three screw. The compressor lost all the pressure. At the end of the Day the system was cleaned the screws of SBN replaced and XB Vacuum pumping. 12/21/2007
CRVThursday, January 3, 2008 09:35:43XB dewar cooldawn started 12/22/2007 . At the end of the Day the system was cold.
crvTuesday, January 29, 2008 17:03:18The system warmed down during this weekend 430mhz dual beam run during the power dips.
crvWednesday, January 30, 2008 14:29:07The system was vac. and is cooling down. 1/30/08
crvWednesday, February 6, 2008 14:07:41Reciver was changed with reciever sbl. The system was originally with compressor ED008 and is now with compressor 009.
crvFriday, April 18, 2008 11:18:33The reciever was giving false reading of the 16 stage temperature changing from 8 to 70. The lakeshore temp. monitor board was replaced.
crvMonday, August 4, 2008 16:32:55dewar temp.monitor problem fixed dewar select board and mux card replaced both where bad.
crvMonday, August 4, 2008 16:33:18system dowm due to maintenance.
GRTuesday, August 12, 2008 16:14:02Cal cables replaced. Calibration pending.
GRTuesday, August 12, 2008 16:15:16Cal cables replaced. Calibration pending.
Stephen Jensensjensen@naic.eduThursday, December 11, 2008 16:33:55The X-band is back on line after having the crosshead replaced.
GRgrivera@naic.eduThursday, February 5, 2009 11:19:06Monday, Nov. 17, 2008: XB crosshead was removed. It has some moisture and a lot of charcoal residue. The cylinder was honed and cleaned. The crosshead was taken to the cryo-shop for cleaning and/or repairing.
GRgrivera@naic.eduThursday, February 5, 2009 11:20:43Tuesday, Dec. 11, 2008: XB crosshead, ED 015 was replaced with ED 007.
crvWednesday, July 1, 2009 14:25:34The recievers cryo pump was replaced on June 1st 2009. for a brand new one.
Stephen Jensensjensen@naic.eduFriday, July 31, 2009 15:02:31The bias box from X-Band was borrowed to use on CBH.
Stephen Jensensjensen@naic.eduMonday, August 3, 2009 16:55:52X-Band is warm after a power failure. Bias box is on CBH.
CRVMonday, January 4, 2010 16:47:20System was removed from Comp. ED015 and connected to ALFA'S new compressor C2 . This compressor has now three cryo pumps (alfa 2, 327MHz , XB).
crioscrios@nbaic.eduMonday, February 22, 2010 14:14:53 System was warmed up on tuesday feb.16th 2010 and started cooldown on Thursday feb.18th 2010. system now connected to ALFA C1.
criosFriday, April 30, 2010 17:26:16The system was disconnected from Compressor C2 and conected on the spare comp. ED015 on April 23rd to allow connecting Alfa second cryo pump to the C1 compressor. This change was made due to C2 comp. failling.
crvThursday, May 6, 2010 14:34:40The system Ch b has an oscilation that was traced to an internal problem of the dewar. The system was turned off today and the dewar will be lowered tomorrow.
crvcrios@naic.eduTuesday, July 20, 2010 10:44:14The dewar was opened and re worked in the last month. The filters connected to the pins of the main dewar connector were removed. Some lose connection on the dewars terminal blocks were retightened. Dewar was close vacuum and thermaly cicled 3 times . No more oscilation were detected. The dewar was installed and cooled on the dome on Friday July 16th 2010.
crvcrios@naic.eduTuesday, December 28, 2010 11:40:05The system was turned off and started warm up yesterday to purge the system and and remove contaminants. Today we purge the system and the cooldown will start after the vac. finishes.
criosWednesday, December 29, 2010 10:48:04vac. pump connected on dewar.
crvcrios@naic.eduMonday, January 3, 2011 15:35:17System started cooldown today at 11:15 am. Nac.was at 1u.
crvMonday, January 17, 2011 16:23:52Guarionex and I checked the dewar and chassis for loose connections and we didn't find none.
crvcrios@naic.eduFriday, May 6, 2011 15:32:50The cooling system was turned off to purge the cryo system bucaesu it was sounding contaminated.
Denis Urbaindurbain@naic.eduMonday, May 9, 2011 11:31:30Start evacuating the cryostat, the vacuum was very poor after the warmup.
Carlos Rioscrios@naic.eduMonday, May 9, 2011 14:55:51System was purged and compressor reconnected and started cool down.
Carlos RiosCrios@naic.eduThursday, February 2, 2012 10:08:40The receiver cooling system was swaped with the SBN and connected with the 327Mhz receiver. This was the XB original setup. January 26th 2012
Carlos Rioscrios@naic.eduThursday, February 2, 2012 10:14:52 The cooling system had a leak and a flexible line on the transmitter room was replaced. After replacing the line the XB and 327 Mhz were contaminated. The system was turned off on Feb. 1st 2012. Started cooldown today.
crioscrios@naic.eduMonday, May 7, 2012 10:13:30The cryo pump sounded contaminated. The system was turned off.
Dana Whitlowdwhitlow@naic.eduWednesday, August 8, 2012 17:43:50Turned XB post-amp chassis back on, so that CALs on other receivers would resume switching modes properly. It is necessary that all post amps be turned on for the CALs to switch properly.
Denis Urbaindurbain@naic.eduFriday, December 7, 2012 16:47:10Small increase in Pol B Tsys, found a loose RF connetion. Tighten the connector and the Tsys is back to normal
Denis Urbaindurbain@naic.eduThursday, June 6, 2013 08:12:39The XB warmed up because the compressor failed. The motor of the fan stopped working and the over temperature sensor was disconected. The phase monitor and other plastic parts had melted because of the excessive temperature. The compressor was repaired, the cold head was cleaned and the vacuum restored. The receiver was cooled again on the 6/4/2013.
Denis Urbaindurbain@naic.eduMonday, July 8, 2013 10:27:25The PolA had some jump in power, moving cable or knocking on the dewar did not affect the output power. The receiver temperature was also higher by a few degrees. Swaping the bias box with the C band hi stop the instability. The bias box of the X band needs to be inspected. The Pol A and Pol B now have similar Tsys. Some cabling problem with pol B are still under investigation.
Felix O. Fernandezfelix.fernandez@naic.eduWednesday, February 19, 2014 07:37:11The XB Rx had been increasing in temperatures since 17-JAN-2014. Finally on 18-JAN-2014 a significant temperature increase was reported. The 16K stage reached 190K. This final temperature spike prompted the decision to warm up the XB Rx. The cold-head was disconnected on 18-JAN-2014.
Felix O. Fernandezfelix.fernandez@naic.eduMonday, March 10, 2014 14:09:48Connected the vacuum pump and cold-head was replaced by cold-head with serial: 17366. The cold-head is currently off and will be turned ON once the satisfactory vacuum is achieved.
Felix O. Fernandezfelix.fernandez@naic.eduTuesday, March 11, 2014 11:38:04Vacuum reached 1.5 mTorr and cooling process was started.
Contamination/Seal Repairfelix.fernandez@naic.eduThursday, May 8, 2014 14:39:38XB had trouble reaching the 17K temp at the second stage. -Warming up (yesterday) -Replaced second stage (thinner one) seal. (Today) -Achieved vacuum of 6 mTorr (Today) -Started cool-down process (Today)
Temperature oscillationsfelix.fernandez@naic.eduWednesday, July 30, 2014 11:13:21The XB 20K stage has been having temperature oscillations for the past several weeks. We believe the problem is related to a poor seal on the second stage o-ring. We will schedule some maintenance time next week to address the issue.
Compressor failurefelix.fernandez@naic.eduWednesday, December 2, 2015 14:54:02Early on Wed Nov-25 the 327/XB compressor failed. After replacing a bad fan motor capacitor, the compressor was placed back online on Monday Nov-30.


Back to Top